OpenStudy (rhi):

how do you express 84 as a product of its prime factors ?

OpenStudy (lucaz):

just divide 84 by the prime numbers 2, 3, 5...

OpenStudy (lucaz):

84/2=42 42/2=21 21/3=7 84=2x2x3x7

OpenStudy (aravindg):

Spli the number as 84=2 X 42 Now 42=7 X6 Now split 6 . Thats all!

OpenStudy (jdoe0001):

the "prime" part threw me off for a sec

OpenStudy (jdoe0001):

 \(\bf 84 = 2 \times 2 \times 3 \times 7\) so-called "simplifying"